# The Convergence Frontier: Integrating Machine Learning and High Performance Quantum Computing for Next-Generation Drug Discovery

### 18 Mar 2026 v1

@@ 4.1 FeNNix-Bio1, a Quantum Foundation Model for Drug Design
In that context, some of us introduced FeNNix-Bio1 [10], a family of foundation machine learning models (S and M versions) designed to enable accurate and reactive atomistic simulations of biological systems with unprecedented speed and scalability. While traditional drug discovery has been hampered by the computational bottleneck of characterizing dynamic interactions between proteins and drug candidates, FeNNix-Bio1 bridges this gap by providing quantum-accurate molecular dynamics without the need for manual parameterization. Trained exclusively on the Ignis inhouse proprietary dataset [10] (see sec. 4.2) — a massive initial collection of over 2.2 million synthetic quantum chemistry configurations — the model employs a novel RaSTER (Range-Separated Transformer with Equivariant Representations) architecture [10]. This architecture utilizes a multi-resolution descriptor that resolves short-range covalent structures with an equivariant Transformer while addressing long-range interactions (up to 11 Å) via a physics-inspired radial basis. To handle the complex charge states ubiquitous in biology, a multi-Qeq embedding was introduced. By non-locally distributing the total charge, it allows the model to differentiate between various charge states in condensed-phase environments. The model’s versatility was demonstrated across a wide spectrum of applications, including the precise prediction of hydration free energies (HFEs) for the Freesolv [128] dataset, where it significantly outperformed traditional GAFF force fields (see ref. [111] and references therein). Furthermore, FeNNix-Bio1 successfully modeled complex phenomena such as the reversible folding of the Chignolin protein—a first for a foundation ML model—and preserved stable binding modes for protein-ligand complexes like Benzamidine-Trypsin. The model also natively handles chemical reactivity, accurately sampling reactive pathways for the conversion of butadiene to cyclobutene through active-learning-driven finetuning. From a performance standpoint, FeNNix-Bio1 is optimized for massively parallel GPU-accelerated systems, showing inference speeds over an order of magnitude faster than existing models like MACE-OFF [127] while maintaining comparable or superior accuracy. By accounting for nuclear quantum effects (NQEs) through an adaptive Quantum Thermal Bath [129, 130] and providing robust uncertainty quantification via the DPOSE method [131], FeNNix-Bio1 establishes a new standard for predictive pharmacological sciences, capable of simulating efficiently systems as large as the SARS-CoV-2 Spike protein with over 1.6 million atoms while providing quantum accurate properties.
@@ 5.2.1 Lambda-ABF-OPES: Maximizing Convergence and Computational Efficiency
PWWP1 Charge-Changing Dual-LAO 24 nsa 0.43 0.92 30x
@@
BRD4 Buried Water Dual-LAO 24 nsa 0.60 0.75 15x
@@
P38 Fragment-like Dual-LAO 24 nsa 0.66 0.82 15x
@@
CHK1 Scaffold-Hopping Dual-LAO 24 nsa 0.52 0.57 27x
@@ 5.3.1 End-to-end quadratic speedups for equilibrium sampling and observable estimation
Thus, quantum algorithms provide compelling asymptotic evidence for accelerated Gibbs sampling, via both quadratic precision improvements and gap amplification. Achieving genuine practical speedups hinges on overcoming the initial state preparation barrier and integrating these primitives into resource-efficient, fault-tolerant architectures [198, 80, 11, 199]. Resolving these challenges would bring quantum advantage in statistical mechanics and its applications significantly closer to reality. In other words, quantum algorithms promise to accelerate drug discovery by facilitating binding free energy estimations and the study of protein-ligand complexes, which Section 4 explains are central to the field. Moreover, some of us have shown that current quantum devices such as Quantinuum’s H2 and 98-qubits Helios can already achieve small quantum Markov Chain Monte Carlo simulations [12], opening a new route for quantum-accelerated atomistic molecular simulations. Let us detailed these recent experiments.

### 19 Mar 2026 v2

@@ 4.1 FeNNix-Bio1, a Quantum Foundation Model for Drug Design
In that context, some of us introduced FeNNix-Bio1 [10], a family of foundation machine learning models (S and M versions) designed to enable accurate and reactive atomistic simulations of biological systems with unprecedented speed and scalability. While traditional drug discovery has been hampered by the computational bottleneck of characterizing dynamic interactions between proteins and drug candidates, FeNNix-Bio1 bridges this gap by providing quantum-accurate molecular dynamics without the need for manual parameterization. Trained exclusively on the Ignis inhouse proprietary dataset [10] (see sec. 4.2) — a massive initial collection of over 2.2 million synthetic quantum chemistry configurations — the model employs a novel RaSTER (Range-Separated Transformer with Equivariant Representations) architecture [10]. This architecture utilizes a multi-resolution descriptor that resolves short-range covalent structures with an equivariant Transformer while addressing long-range interactions (up to 11 Å) via a physics-inspired radial basis.
To handle the complex charge states ubiquitous in biology, a multi-Qeq embedding was introduced. By non-locally distributing the total charge, it allows the model to differentiate between various charge states in condensed-phase environments. The model’s versatility was demonstrated across a wide spectrum of applications, including the precise prediction of hydration free energies (HFEs) for the Freesolv [128] dataset, where it significantly outperformed traditional GAFF force fields (see ref. [111] and references therein). Furthermore, FeNNix-Bio1 successfully modeled complex phenomena such as the reversible folding of the Chignolin protein—a first for a foundation ML model—and preserved stable binding modes for protein-ligand complexes like Benzamidine-Trypsin. The model also natively handles chemical reactivity, accurately sampling reactive pathways for the conversion of butadiene to cyclobutene through active-learning-driven finetuning. From a performance standpoint, FeNNix-Bio1 is optimized for massively parallel GPU-accelerated systems, showing inference speeds over an order of magnitude faster than existing models like MACE-OFF [127] while maintaining comparable or superior accuracy. By accounting for nuclear quantum effects (NQEs) through an adaptive Quantum Thermal Bath [129, 130] and providing robust uncertainty quantification via the DPOSE method [131], FeNNix-Bio1 establishes a new standard for predictive pharmacological sciences, capable of simulating efficiently systems as large as the SARS-CoV-2 Spike protein with over 1.6 million atoms while providing quantum accurate properties.
@@ 5.2.1 Lambda-ABF-OPES: Maximizing Convergence and Computational Efficiency
PWWP1 Charge-Changing Dual-LAO 24 nsa 0.46 0.62 30x
@@
BRD4 Buried Water Dual-LAO 24 nsa 0.64 0.76 15x
@@
P38 Fragment-like Dual-LAO 24 nsa 0.82 0.77 15x
@@
CHK1 Scaffold-Hopping Dual-LAO 24 nsa 0.58 0.48 27x
@@ 5.3.1 End-to-end quadratic speedups for equilibrium sampling and observable estimation
Thus, quantum algorithms provide compelling asymptotic evidence for accelerated Gibbs sampling, via both quadratic precision improvements and gap amplification. Achieving genuine practical speedups hinges on overcoming the initial state preparation barrier and integrating these primitives into resource-efficient, fault-tolerant architectures [198, 80, 11, 199]. Resolving these challenges would bring quantum advantage in statistical mechanics and its applications significantly closer to reality. In other words, quantum algorithms promise to accelerate drug discovery by facilitating binding free energy estimations and the study of protein-ligand complexes, which Section 4 explains are central to the field. Moreover, some of us have shown that current quantum devices such as Quantinuum’s H2 and 98-qubit Helios can already achieve small quantum Markov Chain Monte Carlo simulations [12], opening a new route for quantum-accelerated atomistic molecular simulations. Let us detailed these recent experiments.